how can i make a make a circlegraph to display each set of data.and this hw is due tommorrow so HELP!!!!!!!!!
Answer by xcentaur(357)

It would be more useful to post the data you are required to graph...

But anyways,
first total up everything ur spsd t graph
then find the ratio of each quantity and multiply it by 360 to get the degrees


Wait,let me give an example:
suppose you have 70 apples,50 oranges,30 watermelons,30 bananas.Draw a circle graph of your fruit basket.
Total no of fruits:70+50+30+30=180
Now,angle for apples:(70/180)*360=(70*2)=140 degrees
So draw a circle,and from the centre draw an angle of 140 degrees.shade it and mark that apples.

next oranges,degree:(50/180)*360=50*2=100 degrees
next to the orange part of the circle,draw an angle of 100 degrees.Shade it,and mark that oranges

Do the same for the other two:(30/180*360)=(30*2)=60 degrees


This is how you can draw a circlegraph


Hope this helps,
-xC